# Measurement-Induced Entanglement Phase Transition in Free Fermion Chains

### v1

@@ Abstract
Measurement-induced entanglement phase transitions (MIET) highlight how local measurements drive quantum systems between area-law and volume-law entangled states. This review surveys MIET in free fermion models, focusing on how unitary hopping competes with measurement-induced non-unitarity. We discuss controversies regarding the existence of measurement-induced entanglement phase transition in one dimension, the impact of non-Hermitian skin effects, and potential experimental platforms. We conclude with open challenges, including feedback control and higher-dimensional extensions.
@@ 1 Introduction
Understanding how quantum entanglement evolves under the influence of measurements has emerged as a central theme in contemporary quantum many-body physics. MIETs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12], wherein the entanglement entropy of a quantum system exhibits a sharp transition as a function of the measurement strength or rate, offers a compelling example of how quantum information and non-unitary dynamics can interplay to yield new dynamical phase transitions.
@@ 2.1 From Projective to Weak Measurements
Consider projective measurements where the observable $O$ is expressed in its spectral decomposition as $O=\sum_{y}y\Pi_{y}$. We define a set of measurement operators $\{M_{i}\}$ that satisfy the completeness relation, i.e., $\sum_{i}M_{i}^{\dagger}M_{i}=I$. In the case of projective measurements, these operators are given by the orthogonal projection operators $\Pi_{i}$. Thus, when a measurement yields the outcome $i$, the state of the system is collapsed to the corresponding eigenstate associated with $\Pi_{i}$.
@@
$M(x)=\left(\frac{1}{2\pi\sigma^{2}}\right)^{1/4}\exp{\left[-\frac{(x-\gamma y)^{2}}{4\sigma^{2}}\right]},$ (1)
@@
|$\ket{\psi(x)}=\frac{M(x)\ket{\psi}}{\sqrt{p(x)}}.$ (2)
Due to the coupling $\gamma$ being weak or the pointer width $\sigma$ being large, the exponential factor in the Kraus operator $\exp{[-(x-\gamma y)^{2}/4\sigma^{2}]}$ does not strictly distinguish between the different eigenvalues $i$. As a result, each measurement only minimally perturbs the system, and the outcomes appear as a continuous distribution.
In weak measurements, tuning the parameters $\gamma$ and $\sigma$ allows for a smooth transition from weak to strong measurements. First, when the coupling $\gamma$ is very small or the pointer width $\sigma$ is large, one can expand the exponential function to reveal that each measurement only causes a minor correction to the system’s state. Second, in the projective measurement limit, if the coupling strength $\gamma$ increases or the pointer width $\sigma$ approaches zero, the Gaussian function becomes sharply peaked. In fact, it can be written in $\exp\left[-{(x-\gamma y)^{2}}/{4\sigma^{2}}\right]_{\sigma\to 0}\approx\sqrt{2\pi\sigma^{2}}\,\delta(x-\gamma y)$, where only when $x=\gamma y$ remains nonzero, and the condition $M(x)\propto\Pi_{y}$ is satisfied, thereby recovering the situation of projective measurement.
@@ 3 MIET in Monitored Free Fermion Systems
However, from the perspective of physical implementation, the most natural and simplest model for discussing MIET is monitoring free fermion systems. In these systems, the entanglement bath is provided solely by coherent hopping, while local measurements on the free fermions ensure that the monitored dynamical evolution preserves the Gaussian state structure. This guarantees that such a model remains efficiently simulable even for large system sizes. Ref. [18] was first explored in a free fermion Hamiltonian system that preserves $U(1)$-symmetry of the system. In this study, a quasi-particle pair approach for entanglement in integrable systems was proposed based on generalized hydrodynamics. By employing an ansatz for the collapse of quasi-particle pairs and combining it with large-scale numerical simulations, the work quantitatively demonstrated that the volume-law steady-state entanglement structure arising from unitary dynamics can be disrupted by continuous measurements at any strength (rate). However, the understanding of MIET in locally monitored free fermion systems is still far from complete. A natural question arises: does MIET exist in these systems? If it does, what are its critical properties, and which universality class does it belong to? In fact, there are two competing viewpoints regarding the existence of MIET in the thermodynamic limit for such monitoring free fermion systems.
@@
On the one hand, evidence indicates that monitored free fermion 1D systems exhibit a measurement-induced Berezinskii-Kosterlitz-Thouless (BKT) type phase transition. In this context, Ref. [22] developed an $n$-replica Keldysh field theory approach, and its statistical analysis revealed a nontrivial BKT transition driven by measurement degrees of freedom in a system of spinless Dirac fermions. The study further clarified that while entanglement entropy is a hallmark of MIET, correlation functions can also provide similar information. Additionally, Ref. [19] provided numerical evidence showing that continuously measured free fermion systems can undergo a BKT-type phase transition. Their numerical analysis linked the gapless phase with conformal invariance and the logarithmic scaling of entanglement entropy, see panel (a) of Fig. 1. A series of intriguing variations followed. Refs. [20, 23] examined the fate of the steady-state entanglement MIET in monitoring free fermion chains with long-range coupling. They also studied a fermionic Hamiltonian model with long-range hopping and found that systems with long-range coupling exhibit behavior similar to short-range MIET systems, but only when the long-range decay exponent $\alpha>1.5$; otherwise, MIET is absent, see panel (b) of Fig. 1. Furthermore, Ref. [21] investigated whether introducing a disorder localization potential in a monitored free fermion chain would destroy the BKT universality. Their findings indicate that the competition between measurement and disorder leads to rich steady-state entanglement structures, with the BKT transition remaining robust against disorder, see panel (c) of Fig. 1. Refs. [24, 25] focused on the entanglement dynamics in boundary-driven monitored free fermion systems, exploring the impact of continuous measurements on open quantum systems. Their work revealed that measurement plays a dual role in affecting late-time entanglement, inducing critical behavior, and altering quantum transport properties. In addition, studies on driven many-body systems have demonstrated that a high measurement rate can induce local correlations and anomalous transport. Regarding the fate of periodically driven monitored free fermion systems, Ref. [26] performed a renormalization group analysis and numerical validation of the non-Hermitian sine-Gordon model introduced in Ref. [22] under the high measurement rate limit. The results indicate that the BKT phase transition persists in these periodically driven systems, although its occurrence depends on the specific driving protocol. These variations indicate that the BKT universality class found in monitored free fermion 1D systems is robust against long-range hopping, disorder quenches, and periodic driving. Essentially, this robustness emerges from the competition between measurement-induced non-unitary evolution and the system’s intrinsic unitary dynamics. However, it is well known that changing the dimensionality of a system can significantly affect its critical behavior, scaling laws, and renormalization group currents. Ref. [27] explored continuously measured 2D free fermion systems and discovered that the localization-delocalization MIET in these systems can be described by a unified binary framework involving the $SU(R)$ nonlinear sigma model and the $SU(2R)$ non-Hermitian Hubbard model. Moreover, the wavefunctions exhibit conformal invariance at specific measurement strengths, exceeding the predictions of Anderson localization theory. In other words, these findings suggest that high-dimensional MIET in free fermion systems may feature richer universality classes and novel phase transition mechanisms.
@@
On the other hand, recent studies have suggested that MIET in monitored free fermion 1D systems may not exist in the thermodynamic limit. Specifically, Ref. [28] investigated the competition between unitary dynamics and projective measurements in these systems. Their main finding is that the volume-law phase disappears at any measurement rate, leaving no MIET. Numerical results indicate that in the thermodynamic limit, the average steady-state entanglement entropy exhibits a single transition from volume-law to area-law, irrespective of the measurement rate. Although significant finite-size effects in free fermion systems lead to a residual logarithmic scaling, Ref. [29] further developed a replica Keldysh path-integral method. By deriving a nonlinear sigma model as an effective field theory and combining it with numerical results for time density matrix correlation functions, they demonstrated that monitored free fermion systems eventually tend toward a localized area-law in the thermodynamic limit. This supports the conclusion that an entanglement transition corresponding to MIET is absent in the thermodynamic limit.
@@ 4 MIET in non-Hermitian Free Fermion Systems
To address this question, we will focus on non-Hermitian evolution. In measurement quantum systems, it is always possible to post-select the trajectories corresponding to no-click events and consider only the pure non-Hermitian effective evolution, although implementing this in the laboratory poses an exponentially difficult challenge. This demonstrates that the specific form of the non-Hermitian effective Hamiltonian is entirely determined by the measurement operators. Consequently, effective non-Hermitian Hamiltonian constructed with appropriate measurement operators can, in theory, contribute non-unitary dynamics that compete with the intrinsic unitary dynamics, thereby leading to steady-state entanglement phase transitions. It is worth noting that non-Hermitian quantum mechanics has a long history, dating back to studies on disordered models of non-interacting electrons, non-unitary conformal field theories, and systems with parity-time ($\mathcal{PT}$) symmetry [37, 38]. In recent years, non-Hermitian physics has attracted widespread attention for its unique phenomena [15, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76] that extend beyond those observed in Hermitian systems, while also differing from noisy quantum circuits [77, 78, 79]. In this section, we will review the rich landscape of steady-state entanglement in monitoring free fermion systems under pure non-Hermitian effective evolutions,
@@ 4.1 Entanglement Transitions in Measurement-Induced Non-Hermitian Skin Effects
4.2 Further Results in Measurement-Induced Non-Hermitian Evolutions
@@ 6 Outlook
Nevertheless, despite these substantial advances, many fundamental questions remain unresolved, offering exciting opportunities for future research. One of the most pressing unresolved issues in monitored free fermion systems is how to observe MIET in experimental settings, where overcoming the post-selection problem [98, 99, 100, 88] is crucial for implementation.
@@
On the other hand, there are also many unresolved issues in non-Hermitian evolution without measurement randomness. For instance, it is not yet clear what the entanglement dynamics are in a many-body free fermion system that exhibits a single-particle critical non-Hermitian skin effect [50], nor whether its unique spectral structure can introduce new steady-state entanglement behavior or novel universality classes. Another direction of inquiry concerns the fate of symmetry restoration in monitored-induced non-Hermitian free fermion systems, including whether phenomena such as the quantum Mpemba effect [118, 119, 120, 121] persist. Recent discussions on the entanglement dynamics and emergent dynamical phase transitions in quantum feedback-induced skin effects [31, 32, 34, 33, 35] have suggested that quantum feedback may bridge quantum jump dynamics and non-Hermitian dynamics. However, it remains uncertain whether the skin effect is a necessary condition for these dynamical phase transitions or what mechanisms underlie them.
@@ Acknowledgments
We would like to thank Shan-Zhong Li, Zhi Li, Ching Hua Lee, Shuo Liu, Ze-Chuan Liu, Xhek Turkeshi, Xuyang Huang and Yu-Jun Zhao for helpful discussions. X.-J. Yu was supported by the National Natural Science Foundation of China (Grant No.12405034) and a start-up grant from Fuzhou University. J.-X. Zhong was supported the National Natural Science Foundation of China (Grant No. 12374046 and No. 11874316), the National Basic Research Program of China (Grant No. 2015CB921103), and the Program for Changjiang Scholars and Innovative Research Teams in Universities (Grant No. IRT13093).

### v3

# Measurement-Induced Entanglement Phase Transition in Free Fermion Systems

@@ Abstract
Measurement-induced entanglement phase transitions (MIET) highlight how local measurements drive quantum systems between area-law and volume-law entangled states. This review surveys MIET in free fermion models, focusing on how unitary hopping competes with measurement-induced non-unitarity. We discuss controversies regarding the existence of MIET in one dimension, the impact of non-Hermitian skin effects, and potential experimental platforms. We conclude with open challenges, including feedback control and higher-dimensional extensions.
@@ 1 Introduction
Understanding how quantum entanglement evolves under the influence of measurements has emerged as a central theme in contemporary quantum many-body physics. MIETs [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], wherein the entanglement entropy of a quantum system exhibits a sharp transition as a function of the measurement strength or rate, offers a compelling example of how quantum information and non-unitary dynamics can interplay to yield new dynamical phase transitions.
@@ 2.1 From Projective to Weak Measurements
Consider projective measurements of an observable $O=\sum_{y}y\,\Pi_{y}$, with projectors $\{\Pi_{y}\}$. For later convenience we denote the corresponding measurement operators by the same index set, $\{M_{y}\}$, and they satisfy the completeness relation $\sum_{y}M_{y}^{\dagger}M_{y}=I$. In the projective limit one simply has $M_{y}=\Pi_{y}$, so upon obtaining outcome $y$ the state collapses to $\Pi_{y}\ket{\psi}$.
@@
$M_{y}(x)=\left(\frac{1}{2\pi\sigma^{2}}\right)^{1/4}\exp{\left[-\frac{(x-\lambda y)^{2}}{4\sigma^{2}}\right]},$ (1)
@@
|$\ket{\psi(x)}=\frac{M_{y}(x)\ket{\psi}}{\sqrt{p(x)}}.$ (2)
The parameters $\lambda$ and $\sigma$ govern the nature of the measurement:
  • In the weak measurement regime ($\lambda\ll 1$ or $\sigma\gg 1$) the exponential factor varies slowly with $y$, so each measurement only slightly perturbs the system and provides limited information about the eigenvalue $y$.
  • In the strong measurement limit ($\lambda=1$ and $\sigma\to 0$), the Gaussian function approaches a Dirac delta function $\lim_{\sigma\to 0}({1}/\sqrt{2\pi\sigma^{2}})\exp[-{(x-\gamma y)^{2}}/{2\sigma^{2}}]=\delta(x-y)$. Here, the outcome $x$ is sharply peaked around $x=y$, and the measurement operators become proportional to the projectors $M_{y}(x)\sim\delta(x-y)=\Pi_{y}$, thus recovering projective measurements.
By tuning $\lambda$ and $\sigma$, one can smoothly interpolate between weak and strong measurements, with $\lambda$ controlling the effective scale of the measurement outcome and $\sigma$ determining the resolution of the apparatus.
@@ 3 MIET in Monitored Free Fermion Systems
To begin with, we briefly introduce the general properties of monitored free-fermion systems. Consider a free-fermion system whose dynamics are generated by a quadratic Hamiltonian
$H=\sum_{i,j}c_{i}^{\dagger}h_{ij}c_{j}\;=\;\frac{1}{2}\,\mathbf{C}^{\dagger}h\mathbf{C},\qquad\mathbf{C}^{\mathsf{T}}=(c_{1},\ldots,c_{L},c_{1}^{\dagger},\ldots,c_{L}^{\dagger}),$ (12)
where $h$ is a real, antisymmetric Bogoliubov-de-Gennes matrix. Unitary evolution under $H$ preserves Gaussianity: if the initial state $\rho(0)$ is Gaussian, then the time-evolved state $\rho(t)=e^{-iHt}\rho(0)e^{iHt}$ remains Gaussian and is fully determined by its two-point correlation matrix $G_{ij}(t)=\mathrm{Tr}[\rho(t)\,C_{i}^{\dagger}C_{j}]$. Now introduce a set of measurement operators $\{M_{y}\}$ that commute with the Hamiltonian, $[M_{y},H]=0\;\forall\,y$. Typical examples include (i) strong projectors onto eigen-mode occupations $N_{\alpha}=\tilde{c}_{\alpha}^{\dagger}\tilde{c}_{\alpha}$ and (ii) weak Gaussian POVMs, $M_{y}\;=\;(2\pi\sigma^{2})^{-L/2}\,\exp{[-\!\sum_{\alpha}(N_{\alpha}-y_{\alpha})^{2}/4\sigma^{2}]},$ whose width $\sigma$ tunes the measurement strength. Because $M_{y}$ and $H$ share the same eigenbasis, both the post-measurement conditional state $\rho_{y}=M_{y}\rho M_{y}^{\dagger}/\mathrm{Tr}[M_{y}\rho M_{y}^{\dagger}]$ and the averaged state $\sum_{y}M_{y}\rho M_{y}^{\dagger}$ remain Gaussian. At the level of correlations this amounts to an affine map $G\longmapsto G^{\prime}=RGR^{\mathsf{T}}+Q$, with an orthogonal matrix $R$ and symmetric matrix $Q$ fixed solely by $\sigma$. Taking the continuous-monitoring limit one obtains a closed equation ${\mathrm{d}G}/{\mathrm{d}t}=-i\,[h,G]-\sigma^{-2}\delta t(G-DGD),$ where $D$ projects onto mode-occupation space. Hence any measurement that commutes with $H$ preserves the Gaussian structure, ensures Wick’s theorem remains valid, and allows all entanglement and coherence measures to be extracted from $G$ alone, providing a self-contained and efficient framework for analysing MIET in free-fermion systems.
@@
However, from the perspective of physical implementation, the most natural and simplest model for discussing MIET is monitoring free fermion systems. In these systems, the entanglement bath is provided solely by coherent hopping, while local measurements on the free fermions ensure that the monitored dynamical evolution preserves the Gaussian state structure. This guarantees that such a model remains efficiently simulable even for large system sizes. Ref. [24] was the pioneer of consider continuous measurements on the free fermions. Than, Ref. [25] was explored in a free fermion Hamiltonian system that preserves $U(1)$-symmetry of the system. In this study, a quasi-particle pair approach for entanglement in integrable systems was proposed based on generalized hydrodynamics. By employing an ansatz for the collapse of quasi-particle pairs and combining it with large-scale numerical simulations, the work quantitatively demonstrated that the volume-law steady-state entanglement structure arising from unitary dynamics can be disrupted by continuous measurements at any strength (rate). However, the understanding of MIET in locally monitored free fermion systems is still far from complete. A natural question arises: does MIET exist in these systems? If it does, what are its critical properties, and which universality class does it belong to? In fact, there are two competing viewpoints regarding the existence of MIET in the thermodynamic limit for such monitoring free fermion systems.
@@
On the one hand, evidence indicates that monitored free fermion 1D systems exhibit a measurement-induced Berezinskii-Kosterlitz-Thouless (BKT) type phase transition. In this context, Ref. [29] developed an $n$-replica Keldysh field theory approach, and its statistical analysis revealed a nontrivial BKT transition driven by measurement degrees of freedom in a system of spinless Dirac fermions. The study further clarified that while entanglement entropy is a hallmark of MIET, correlation functions can also provide similar information. Additionally, Ref. [26] provided numerical evidence showing that continuously measured free fermion systems can undergo a BKT-type phase transition. Their numerical analysis linked the gapless phase with conformal invariance and the logarithmic scaling of entanglement entropy, see panel (a) of Fig. 1. A series of intriguing variations followed. Refs. [27, 30] examined the fate of the steady-state entanglement MIET in monitoring free fermion chains with long-range coupling. They also studied a fermionic Hamiltonian model with long-range hopping and found that systems with long-range coupling exhibit behavior similar to short-range MIET systems, but only when the long-range decay exponent $\alpha>1.5$; otherwise, MIET is absent, see panel (b) of Fig. 1. Furthermore, Ref. [28] investigated whether introducing a disorder localization potential in a monitored free fermion chain would destroy the BKT universality. Their findings indicate that the competition between measurement and disorder leads to rich steady-state entanglement structures, with the BKT transition remaining robust against disorder, see panel (c) of Fig. 1. Refs. [31, 32] focused on the entanglement dynamics in boundary-driven monitored free fermion systems, exploring the impact of continuous measurements on open quantum systems. Their work revealed that measurement plays a dual role in affecting late-time entanglement, inducing critical behavior, and altering quantum transport properties. In addition, studies on driven many-body systems have demonstrated that a high measurement rate can induce local correlations and anomalous transport. Regarding the fate of periodically driven monitored free fermion systems, Ref. [33] performed a renormalization group analysis and numerical validation of the non-Hermitian sine-Gordon model introduced in Ref. [29] under the high measurement rate limit. The results indicate that the BKT phase transition persists in these periodically driven systems, although its occurrence depends on the specific driving protocol. These variations indicate that the BKT universality class found in monitored free fermion 1D systems is robust against long-range hopping, disorder quenches, and periodic driving. Essentially, this robustness emerges from the competition between measurement-induced non-unitary evolution and the system’s intrinsic unitary dynamics. However, it is well known that changing the dimensionality of a system can significantly affect its critical behavior, scaling laws, and renormalization group currents. Ref. [34] explored continuously measured 2D free fermion systems and discovered that the localization-delocalization MIET in these systems can be described by a unified binary framework involving the $SU(R)$ nonlinear sigma model and the $SU(2R)$ non-Hermitian Hubbard model. Moreover, the wavefunctions exhibit conformal invariance at specific measurement strengths, exceeding the predictions of Anderson localization theory. In other words, these findings suggest that high-dimensional MIET in free fermion systems may feature richer universality classes and novel phase transition mechanisms. Additionally, Ref. [35] presents a classification of topological insulators and superconductors, analogous to the periodic table of topological phases in open quantum systems. It provides a systematic framework for exploring how monitored free fermion systems exhibit topological features across different symmetry classes and spacetime dimensions.
@@
On the other hand, recent studies have suggested that MIET in monitored free fermion 1D systems may not exist in the thermodynamic limit. Specifically, Ref. [36] investigated the competition between unitary dynamics and projective measurements in these systems. Their main finding is that the volume-law phase disappears at any measurement rate, leaving no MIET. Numerical results indicate that in the thermodynamic limit, the average steady-state entanglement entropy exhibits a single transition from volume-law to area-law, irrespective of the measurement rate. Although significant finite-size effects in free fermion systems lead to a residual logarithmic scaling, Ref. [37] further developed a replica Keldysh path-integral method. By deriving a nonlinear sigma model as an effective field theory and combining it with numerical results for time density matrix correlation functions, they demonstrated that monitored free fermion systems eventually tend toward a localized area-law in the thermodynamic limit. This supports the conclusion that an entanglement transition corresponding to MIET is absent in the thermodynamic limit. Ref. [38] presents an exact analytical solution for the effective continuum field theory of a monitored free fermion system with conserved charge, and also demonstrates the absence of MIET in such charge-conserving free fermion systems.
@@ 4 MIET in non-Hermitian Free Fermion Systems
To address this question, we will focus on non-Hermitian evolution. In measurement quantum systems, it is always possible to post-select the trajectories corresponding to no-click events and consider only the pure non-Hermitian effective evolution, although implementing this in the laboratory poses an exponentially difficult challenge. This demonstrates that the specific form of the non-Hermitian effective Hamiltonian is entirely determined by the measurement operators. Consequently, effective non-Hermitian Hamiltonian constructed with appropriate measurement operators can, in theory, contribute non-unitary dynamics that compete with the intrinsic unitary dynamics, thereby leading to steady-state entanglement phase transitions. It is worth noting that non-Hermitian quantum mechanics has a long history, dating back to studies on disordered models of non-interacting electrons, non-unitary conformal field theories, and systems with parity-time ($\mathcal{PT}$) symmetry [46, 47]. In recent years, non-Hermitian physics has attracted widespread attention for its unique phenomena [21, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86] that extend beyond those observed in Hermitian systems, while also differing from noisy quantum circuits [87, 88, 89]. In this section, we will review the rich landscape of steady-state entanglement in monitoring free fermion systems under pure non-Hermitian effective evolutions,
@@ 4.1 Entanglement Transitions in Measurement-Induced Non-Hermitian Skin Effects
4.2 Entanglement Transitions beyond Non-Hermitian Skin Effects
@@ 6 Outlook
Nevertheless, despite these substantial advances, many fundamental questions and controversies remain unresolved, offering exciting avenues for future research. One key issue is whether a true BKT transition exists in monitored free-fermion systems. Numerical simulations, limited by finite system sizes and evolution times, can easily misidentify exponential divergences as power laws, while variations in measurement protocols and replica-continuation schemes modify the marginal operators in the effective field theory, leading some analyses to predict a BKT transition and others to rule it out. Decisive verification therefore demands ultra-large-scale simulations coupled with a rigorous finite-size-scaling analysis of multiple observables and a controlled renormalization gruop (RG) treatment within the replica field theory; the principal obstacles are the sheer computational cost and the theoretical ambiguity that arises when replica continuation and RG flow do not commute. Another pressing challenge is the experimental observation of MIET in free-fermion platforms, where overcoming the post-selection problem [108, 109, 110, 98] is essential for any practical implementation.
@@
On the other hand, there are also many unresolved issues in non-Hermitian evolution without measurement randomness. For instance, it is not yet clear what the entanglement dynamics are in a many-body free fermion system that exhibits a single-particle critical non-Hermitian skin effect [59], nor whether its unique spectral structure can introduce new steady-state entanglement behavior or novel universality classes. Another direction of inquiry concerns the fate of symmetry restoration in monitored-induced non-Hermitian free fermion systems, including whether phenomena such as the quantum Mpemba effect [128, 129, 130, 131, 132, 133] persist. Recent discussions on the entanglement dynamics and emergent dynamical phase transitions in quantum feedback-induced skin effects [40, 41, 43, 42, 44] have suggested that quantum feedback may bridge quantum jump dynamics and non-Hermitian dynamics. However, it remains uncertain whether the skin effect is a necessary condition for these dynamical phase transitions or what mechanisms underlie them.
@@ Acknowledgments
We would like to thank Shan-Zhong Li, Zhi Li, Ching Hua Lee, Shuo Liu, Ze-Chuan Liu, Xhek Turkeshi, Xuyang Huang and Yu-Jun Zhao for helpful discussions. X.-J. Yu was supported by the National Natural Science Foundation of China (Grant No.12405034) and a start-up grant from Fuzhou University. J.-X. Zhong was supported the National Natural Science Foundation of China (Grant No.12374046 and No.11874316), the Shanghai Science and Technology Innovation Action Plan (Grant No.24LZ1400800), the National Basic Research Program of China (Grant No.2015CB921103), and the Program for Changjiang Scholars and Innovative Research Teams in Universities (Grant No.IRT13093).